# Graph-based Confidence Calibration for Large Language Models

### v1

@@ Abstract
One important approach to improving the reliability of large language models (LLMs) is to provide accurate confidence estimations regarding the correctness of their answers. However, developing a well-calibrated confidence estimation model is challenging, as mistakes made by LLMs can be difficult to detect. We propose a novel method combining the LLM’s self-consistency with labeled data and training an auxiliary model to estimate the correctness of its responses to questions. This auxiliary model predicts the correctness of responses based solely on their consistent information. To set up the learning problem, we use a weighted graph to represent the consistency among the LLM’s multiple responses to a question. Correctness labels are assigned to these responses based on their similarity to the correct answer. We then train a graph neural network to estimate the probability of correct responses. Experiments demonstrate that the proposed approach substantially outperforms several of the most recent methods in confidence calibration across multiple widely adopted benchmark datasets. Furthermore, the proposed approach significantly improves the generalization capability of confidence calibration on out-of-domain (OOD) data.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities across various natural language processing tasks such as question answering (Wei et al., 2022; Shen et al., 2023; Zheng et al., 2023; Qin et al., 2023; Singhal et al., 2023), text summarization (Tang et al., 2023; Deroy et al., 2023; Tam et al., 2023; Roit et al., 2023), and even creative writing (Gómez-Rodríguez & Williams, 2023; Wang et al., 2024; Deng et al., 2024). Despite their impressive performance, LLMs often give wrong answers in question-answering tasks. There is an urgent need to check the correctness of LLMs’ responses. One particularly interesting question is to calibrate the confidence levels of the correctness of responses from LLMs (Kuhn et al., 2022; Ulmer et al., 2022; He & Jiang, 2023; Van Landeghem et al., 2022; Vazhentsev et al., 2023; Ulmer et al., 2024). Accurate confidence estimation is vital for deploying LLMs in the real world, as it allows users to gauge the reliability of the model’s predictions and make informed decisions based on these outputs. On the contrary, miscalibrated confidence may lead to over-trust in incorrect responses or doubts about the correct ones. For example, a misleading response may steer a patient in the wrong direction when making health decisions; it may also lead an investor to impulsive investment decisions.
@@
In this work, we consider calibrating the confidence with the correctness of LLMs’ responses. This task is challenging in several aspects. First, due to LLMs’ superior ability to generate text, mistakes in an LLM’s response are usually at the semantic level, making it hard to detect even for humans. There are methods using an auxiliary Language Model (e.g., DeBERTa (He et al., 2020)) to check whether the LLM’s response answers the question Ulmer et al. (2024). Since the LLM is supposed to be much stronger than the LM, the LLM should be able to avoid most mistakes that can be detected by an LM; this type of method may omit a significant fraction of wrong answers. Second, it is hard to detect mistakes from the LLM’s internal working mechanism. Because the LLM uses many hidden layers to process the information, it is hard to discern the signal from a small number of hidden units. Even if this is possible, it is not easy to apply this type of method to black-box LLMs.
Recently, there has been some progress in quantifying the model’s own confidence in a response through consistency among the model’s responses (Chen & Mueller, 2023; Lin et al., 2024). If the LLM always gives similar responses, then there is less uncertainty, and any one of these responses tends to have a high probability of being correct. In particular, the results show that the model’s own confidence in its response has a strong correlation with the correctness of the response. This new development leads to an important research question: whether we can calibrate the confidence of correctness from consistency among the LLM’s responses.
In this work, we develop a new method to calibrate the confidence of correctness from the consistency of an LLM’s responses. To achieve this goal, we train a separate calibration model to predict the correctness of the LLM’s responses. To get the input to the calibration model, we form a similarity graph over the LLM’s multiple responses to the same question. The similarity graph encodes information about the consistency between LLM’s responses. A response consistent with more answers tends to have a higher likelihood of being correct, so the consistency graph is predictive of the responses’ correctness. In our new approach, the calibration model only considers the consistency among responses without processing any actual language information. Thus, we can use a relatively simple and efficient model.
Our model achieves premium performance in the empirical study. Compared with previous calibration methods, our model has much better calibration performance because of the usage of consistency graphs. Compared to prior methods based on consistency inputs, our method improves not only the calibration performance but also the ranking performance. Furthermore, our method improves the generalizability of confidence calibration for out-of-domain settings, demonstrating the advantage of using a separate learning model. In summary, our main contributions are:
  • Graph-based confidence calibration method: We propose a novel graph-based confidence calibration approach to improve the reliability of LLMs.
@@
  • Enhanced calibration performance: Our evaluations demonstrate that the proposed method substantially outperforms recent methods in confidence calibration across several widely used benchmark datasets.
  • Improved OOD generalizability: Evaluations on OOD confidence calibration show that our graph-based approach significantly improves generalizability in OOD settings.
@@ 2 Related Work
Due to the urgent need to improve the reliability of LLMs, confidence estimation and calibration for these models have become active areas of research. Existing research in LLM uncertainty quantification can be summarized into two main categories: uncertainty quantification and confidence calibration (Geng et al., 2023). Confidence estimation for short responses (e.g., for multi-choice or yes-no questions) is generally less complicated than for long responses (Ye et al., 2024). For a brief response, the LLM’s output logits are informative about its confidence; the easy comparisons of responses to the true answer facilitate both calibration and evaluation. Confidence estimation for long responses cannot simply depend on LLM’s output logits (Duan et al., 2023; Bakman et al., 2024) because the logits indicate more about the probability of text and less about the semantics behind it. There are also methods using the internal state of an LLM (Ren et al., 2022), but it is not always available to have such information about the LLM interface.
@@
Another approach is to check the LLM’s consistency in its responses. Kotelanski et al. (2023) demonstrate that repeated sampling and consistency checks across multiple outputs can serve as reliable proxies for model confidence. Manakul et al. (2023) generates multiple responses from the LLM and checks the consistency between responses using various methods, including querying the LLM. Chen & Mueller (2023) combines the consistency between responses and the LLM’s self-reflection certainty to quantify the uncertainty. Kuhn et al. (2022) considers confidence from semantic equivalence and proposes a method based on clustering of responses. Lin et al. (2024) organize responses in a graph with their pairwise semantic similarity and then extract graph statistics for confidence estimation. Zhang et al. (2024) examines methods of comparing responses via entailment and contradiction relationships. These studies highlight the importance of semantic consistency in confidence estimation. These methods are evaluated by comparing their estimated confidence values against the correctness of responses. The correct labels are often obtained by reference matching (Huang et al., 2024), checking whether responses match true answers with a particular similarity measure.
@@
To better calibrate the confidence estimation, some methods directly use correctness labels in their calibration procedures. Mielke et al. (2022) trains a calibrator to predict the correctness of a response for a given question. With a similar idea, Ulmer et al. (2024) trains a language model (e.g., DeBERTa) based on question-response pairs to predict the probability of responses’ correctness. Based on SelfCheckGPT (Manakul et al., 2023) and JAFC (Tian et al., 2023), Chen et al. (2024) train supervised models to reduce grouping losses and improve the confidence estimation. The method by Liu et al. (2024) uses an LLM’s latent representations to predict the correctness of responses. Detommaso et al. (2024) uses the “multicalibration” technique to calibrate the probability of correctness. (Fadeeva et al., 2023) offers a detailed comparative study of various confidence estimation methods, providing empirical evidence on their effectiveness across different tasks. However, these studies have not sufficiently exploited response consistency to predict the probabilities of the responses being correct.
@@ 3 Method
Background:  The formulation of semantic equivalence (Kuhn et al., 2022) provides a framework for our analysis. Let $\mathcal{R}$ be the space of all possible responses. Given a question $q$, the space $\mathcal{R}$ is divided into a set $\mathcal{C}_{q}$ of semantic classes: $\mathcal{R}=\cup_{C\in\mathcal{C}_{q}}C$ and $C^{\prime}\cap C=\emptyset$ for any two different semantic classes $C,C^{\prime}\in\mathcal{C}_{q}$. For two responses $r_{1},r_{2}\in C$ in the same equivalent class, they are considered as the same semantic response: if one is the correct answer, the other is correct as well, and vice versa. Then, we can consider the quality of the LLM’s responses at the semantic level. In particular, a semantic response $C$ has probability
@@
Here $p(r|q)$ is the probability of a single response from the LLM.
@@
We assume that each cluster cluster $\tilde{C}$ is from a different semantic class $C$, then $p(C|q)$ can be approximated by
@@
Now, we depart from the setup of semantic classes and consider the correctness of responses. Let $C^{*}$ be the correct semantic answer to question $q$. Without knowing which responses are correct answers, a common assumption is that the model’s confidence reflects the correctness, that is, the model’s confidence in a semantic response is approximately the probability of correctness, then
@@
It says that the more certain the model is about a semantic response, the more likely the response is correct. Conversely, a wide variation in the LLM’s responses indicates low confidence in all responses $r_{i}$ and low accuracy. This pattern is also found in previous studies (Kuhn et al., 2022; Lin et al., 2024).
While we agree that a positive correlation exists between the LLM’s confidence and the probability of correctness, we do not believe that they are equal, as shown in equation 3. Therefore, we need further calibration to reflect the probability of correctness.
Figure 1: The overall framework. Given an input question, our approach first samples multiple responses from the LLM and then constructs a similarity graph based on the samples, which serves as the input for the GNN model to calibrate the confidence of LLM responses. The left part is an example of question and responses from LLM (“R” denotes “response,” “C” denotes “confidence”).
@@ 3.1 Confidence calibration as graph learning problem
Now, we set a supervised learning problem and train a model to calibrate the confidence of the correctness of responses. We first consider the correctness labels of the LLM’s responses. In the supervised setting, we have a correct answer $r^{*}$ to the question $q$. Then $r^{*}$ to assign correctness labels to sampled responses $\{r_{1},r_{2},...,r_{n}\}$ for the same question $q$. In our work, we use the ROUGE similarity. Specifically, we compute the ROUGE similarity $\mathrm{sim}_{R}(r_{i},r^{*})$ between a sampled response and the correct answer to decide the correctness label.
@@
Here $\mathbbm{1}[\cdot]$ is one if the condition is true or 0 otherwise. The ROUGE metric is reasonably accurate in measuring semantic similarity between short sentences (Lin & Och, 2004).
@@
Then, we treat the problem as a node classification problem (Xiao et al., 2022). In particular, we run a Graph Neural Network (GNN) $\mathrm{gnn}(\cdot)$ to predict the probability of each response being correct
@@
To provide clustering information to the GNN, we first run the K-means clustering algorithm on the responses’ embeddings and assign cluster IDs from 0 to $K-1$ based on the order from largest to smallest (ties are randomly broken). Then, we feed each response’s cluster membership as a one-hot feature input to the GNN. Therefore, the GNN’s predictions are purely based on the relationships between responses in semantic space. We choose NOT to feed in the embedding vectors of responses to avoid the GNN’s dependency on textual information. This helps the GNN generalize questions from different domains. The overall framework is shown in Fig 1.
@@
$\ell_{q}=-\sum_{i=1}^{n}y_{i}\tilde{p}_{i}+(1-y_{i})\log(1-\tilde{p}_{i})$ (7)
Note that the loss is consistent marginally since the loss is minimized when $\tilde{p}_{i}=p(y_{i}|G)$. An alternative approach is to minimize the squared error $(y_{i}-\tilde{p}_{i})$, from which we get similar performances, so we choose the cross-entropy loss. A further consideration is to explicitly consider the similarity between $\tilde{p}_{i}$ and $\tilde{p}_{j}$ given the response similarity $w_{ij}$. We leave such exploration to the future.
@@ 3.2 Improve the estimation through multiple prompts
It is well known that the syntactic form of a question influences responses and introduces additional variance. To reduce this variance and evaluate the LLM’s semantic consistency, we analyze the LLM’s responses to multiple prompts derived from the same question. These responses are treated as answers to the same semantic question. We then apply the same method as before to predict the correctness of each response.
In particular, we rephrase the original question $q$ into $k$ different forms $\{q_{1},...,q_{k}\}$ while maintaining the original sentence’s semantic meaning. We employ a multiple rephrased questions strategy for answer sampling. Specifically, we prompt the GPT-4 to give $k$ different but with the same meaning rephrased questions for the given question $q$. Then, we sample $n/k$ responses from the LLM for each rephrased question and still get a total of $n$ responses, from which the confidence calibration is the same as we have described above. The model is more likely to produce diverse responses for questions about which the LLM is less certain. Confidence calibration is more accurate in this scenario because the model’s uncertainty becomes more apparent.

### v2

@@ Abstract
Reliable confidence estimation is essential for enhancing the trustworthiness of large language models (LLMs), especially in high-stakes scenarios. Despite its importance, accurately estimating confidence in LLM responses remains a significant challenge. In this work, we propose using an auxiliary learning model to assess response correctness based on the self-consistency of multiple outputs generated by the LLM. Our method builds a consistency graph to represent the agreement among multiple responses and uses a graph neural network (GNN) to estimate the likelihood that each response is correct. Experiments demonstrate that this method has strong calibration performance on various benchmark datasets and generalizes well to out-of-domain cases.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities across various natural language processing tasks such as question answering (Wei et al., 2022; Shen et al., 2023; Zheng et al., 2023; Qin et al., 2023; Singhal et al., 2023), text summarization (Tang et al., 2023; Deroy et al., 2023; Tam et al., 2023; Roit et al., 2023), and even creative writing (Gómez-Rodríguez & Williams, 2023; Wang et al., 2024; Deng et al., 2024). Despite their impressive performance, LLMs often give incorrect responses in question-answering tasks. One particularly important challenge lies in calibrating the confidence levels of LLM-generated responses (Kuhn et al., 2022; Ulmer et al., 2022; Van Landeghem et al., 2022; Vazhentsev et al., 2023; Ulmer et al., 2024). Accurate confidence estimation is vital for deploying LLMs in the real world, as it enables users to gauge the reliability of the model’s predictions and make informed decisions accordingly. On the contrary, miscalibrated confidence may lead to over-reliance on incorrect responses or unnecessary skepticism toward the correct ones. For example, a misleading response may steer a patient in the harmful direction when making health decisions; it may also cause an investor to make impulsive financial choices.
@@
In this work, we focus on calibrating LLMs’ confidence to better reflect the correctness of their responses. This task is challenging in several aspects. First, due to LLMs’ superior ability to generate text, mistakes in their response often occur at the semantic level, making them hard to detect even for humans. There are methods using an auxiliary Language Model (e.g., DeBERTa (He et al., 2020)) to verify whether the LLM’s response appropriately answers the question (Ulmer et al., 2024). Since the LLM is supposed to be much stronger than the LM, the LLM should be able to avoid most mistakes that can be detected by an LM; so this type of method may omit a significant fraction of wrong answers. Second, it is hard to detect mistakes from the LLM’s internal working mechanism. Because the LLM uses many hidden layers to process the information, it is hard to discern the signal from a small number of hidden units. Even if this is possible, it is not easy to apply this type of method to black-box LLMs.
Recently, there has been some progress in quantifying the model’s confidence in its own responses through consistency among the outputs generated by the model itself (Chen & Mueller, 2023; Lin et al., 2024). These approaches demonstrate a strong correlation between an LLM’s self-consistency and the actual correctness of its responses. However, because these methods depend on hand-crafted features, they often fail to accurately calibrate confidence, resulting in a mismatch between predicted confidence levels and the true accuracy of the answers. This raises an important research question: can we improve confidence calibration by learning from patterns of consistency across the LLM’s own responses?
In this work, we propose an auxiliary learning model to improve confidence calibration. We begin by constructing similarity graphs from the LLM’s multiple responses to the same questions, where graph edges represent the degree of agreement between responses. We then train a separate calibration model using graph neural networks (GNNs) to predict the correctness of each response. The key insight is that consistency among responses carries strong signals of correctness – for instance, a response that aligns well with many others is more likely to be accurate. Importantly, our model operates solely on response consistency and does not analyze the actual language content. This work focuses on practical scenarios where correctness reflects alignment with the training data, and does not address the case where training data itself contains consistent but incorrect information. The latter case is a more challenging scenario investigated by ongoing research (Biester et al., 2024; Shi et al., 2023; Krishnan & Wu, 2019).
We further investigate the problem of transferring a calibration model across different question domains, which is crucial when target domains lack sufficient training data. Despite its importance, this problem has received limited attention in the literature. Our study demonstrates that the proposed auxiliary calibration model can generalize to new domains with minimal performance degradation. This generalization is from the observation that self-consistency can serve as a broadly applicable signal for confidence, enabling the learning model that relies solely on self-consistency to achieve strong transfer performance.
We evaluate the performance of the proposed method with an extensive empirical study that includes four datasets from different question domains. Empirical results show that our method achieves strong performance. Besides the improved calibration performance, our model enhances the ranking of an LLM’s responses. The study has also tested our model and competing models in out-of-domain settings. The results show that the proposed method shows robust performance when generalizing to new domains.
In summary, our main contributions are:
  • A learning-based GNN framework: We propose a learning-based framework leveraging GNNs to calibrate confidence values of LLMs’ responses.
@@
  • Enhanced calibration performance: We conduct an extensive empirical study to evaluate the proposed method and show that it substantially outperforms recent methods in confidence calibration across several widely used benchmark datasets.
  • Improved out-of-domain generalizability: We investigate the scenario of out-of-domain (OOD) confidence calibration and show the superior performance of the proposed method in this setting.
@@ 2 Related Work
Due to the urgent need to improve the reliability of LLMs, confidence estimation and calibration for these models have become active areas of research. Existing research in LLM uncertainty quantification can be summarized into two main categories: uncertainty quantification and confidence calibration (Geng et al., 2023). Confidence estimation for short responses (e.g., for multi-choice or yes-no questions) is generally less complicated than for long responses (Ye et al., 2024). For a brief response, the LLM’s output logits are informative about its confidence; the easy comparisons of responses to the true answer facilitate both calibration and evaluation. Confidence estimation for long responses cannot simply depend on LLM’s output logits (Duan et al., 2023; Bakman et al., 2024) because the logits indicate more about the probability of text and less about the semantics behind it. There are also methods using the internal state of an LLM (Ren et al., 2022; Beigi et al., 2024), but it is not always available to have such information about the LLM interface.
@@
Another approach is to check the LLM’s consistency in its responses. Kotelanski et al. (2023) demonstrate that repeated sampling and consistency checks across multiple outputs can serve as reliable proxies for model confidence. Manakul et al. (2023) generate multiple responses from the LLM and check the consistency between responses using various methods, including querying the LLM. Chen & Mueller (2023) combine the consistency between responses and the LLM’s self-reflection certainty to quantify the uncertainty. Kuhn et al. (2022) consider confidence from semantic equivalence and proposes a method based on clustering of responses. Lin et al. (2024) organize responses in a graph with their pairwise semantic similarity and then extract graph statistics for confidence estimation. Zhang et al. (2024) examine methods of comparing responses via entailment and contradiction relationships. These studies highlight the importance of semantic consistency in ranking an LLM’s responses. However, manually designed features are limited in their ability to capture the full extent of self-consistency among LLM responses, leading to poor calibration performance.
@@
To better calibrate the confidence estimation, some methods directly use correctness labels in their calibration procedures. Mielke et al. (2022) train a calibrator to predict the correctness of a response for a given question. With a similar idea, Ulmer et al. (2024) train a language model (e.g., DeBERTa) based on question-response pairs to predict the probability of responses’ correctness. Based on SelfCheckGPT (Manakul et al., 2023) and JAFC (Tian et al., 2023), Chen et al. (2024) train supervised models to reduce grouping losses and improve the confidence estimation. The method by Liu et al. (2024) uses an LLM’s latent representations to predict the correctness of responses. Detommaso et al. (2024) use the “multicalibration” technique to calibrate the probability of correctness. Fadeeva et al. (2023) offer a detailed comparative study of various confidence estimation methods, providing empirical evidence on their effectiveness across different tasks. However, these studies have not sufficiently exploited response consistency to predict the probabilities of the responses being correct.
@@ 3 Method
Background:  The formulation of semantic equivalence (Kuhn et al., 2022) provides a framework for our analysis. Let $\mathcal{R}$ be the space of all possible responses. Given a question $q$, the space $\mathcal{R}$ is divided into a set $\mathcal{C}_{q}$ of semantic classes: $\mathcal{R}=\cup_{C\in\mathcal{C}_{q}}C$ and $C^{\prime}\cap C=\emptyset$ for any two different semantic classes $C,C^{\prime}\in\mathcal{C}_{q}$. Two responses $r_{1},r_{2}\in C$ in the same equivalent class are considered as the same semantic response: if one is the correct answer, the other is correct as well. Under an LLM, the semantic response $C$ has probability:
@@
Here $p(r|q)$ is the probability of a single response from the LLM. We also say that $p(C|q)$ is the LLM’s confidence in the semantic response $C$.
@@
We assume that each cluster $\tilde{C}$ is from a different semantic class $C$, then $p(C|q)$ can be approximated by
@@
Now, we depart from the setup of semantic classes and consider the correctness of responses. Let $C^{*}$ be the correct semantic answer to question $q$. Without knowing $C^{*}$, a common assumption is that the model’s confidence $p(C|q)$ in the semantic response $C$ reflects the correctness. For example, $p(C|q)$ is approximately the probability of correctness:
@@
With this assumption, the more certain the model is about a semantic response, the more likely the response is correct. Conversely, a wide variation in the LLM’s responses indicates low confidence in all responses $r_{i}$ and low accuracy. This pattern is also found in previous studies (Kuhn et al., 2022; Lin et al., 2024).
While a positive correlation exists between the LLM’s confidence and the probability of correctness, the two quantities are unlikely to have the equal relationship shown in (3). Therefore, we need further calibration to estimate the probability of correctness.
Figure 1: The overall framework of our confidence calibration model. Given an input question, our approach first generates multiple responses from the LLM and constructs a similarity-weighted graph based on these responses. This graph serves as the input for the GNN model, which calibrates the confidence of the LLM responses. In the weighted graph, the edge weight $w_{ij}$ is defined as $\mathrm{sim_{cos}}(\mathrm{emb}(r_{i}),\mathrm{emb}(r_{j}))$, where ${}~{}i,j=1,\ldots,n.$ A higher weight indicates greater similarity between the responses. We also use cluster memberships as the node features to enhance the performance.
@@ 3.1 Confidence calibration as graph learning problem
Now, we set a supervised learning problem and train a model to calibrate the confidence of the correctness of responses. We first consider the correctness labels of the LLM’s responses. In the supervised setting, we have a ground-truth answer $a$ to the question $q$. Then we use $a$ to assign correctness labels to sampled responses $\{r_{1},r_{2},...,r_{n}\}$ for the same question $q$.
In our work, we have two approaches to label these responses. In the first approach, we use the ROUGE similarity. Specifically, we compute the ROUGE similarity $\mathrm{sim}_{R}(r_{i},a)$ between a sampled response and the correct answer to decide the correctness label.
@@
Here $\mathbbm{1}[\cdot]$ is one if the condition is true or 0 otherwise. As shown in previous studies (Lin & Och, 2004) and our own study, the ROUGE metric is reasonably accurate in measuring semantic similarity between short sentences. We follow the previous work, and set $\tau=0.3$ (Kuhn et al., 2022).
In the second method, we utilize the LLM to generate correctness labels. Specifically, we provide the question $q$ and the standard answer $a$ as the context, then ask whether the response $r_{i}$ answers the question $q$. The response from the LLM is then used as the label for $r_{i}$. We denote the procedure as
$y_{i}=\mathrm{llm_{y}}(q,a,r_{i})$ (5)
We provide the prompt for labeling in the Appendix F.
Both methods are automatic and can scale up to large datasets. To guaranteed the labeling quality, we also include a relatively small set of manual labels in our study. For each question, a human labeler inspects the question $q$ and the true answer $a$, and then assign the correctness label $y_{i}$ to a response $r_{i}$.
@@
Then, we treat the problem as a node classification problem (Xiao et al., 2022). In particular, we run a GNN $\mathrm{gnn}(\cdot)$ to predict the probability of each response being correct
@@
To provide clustering information to the GNN, we first run the K-means clustering algorithm on the responses’ embeddings and assign cluster ids from 0 to $K-1$ based on the order from largest to smallest (ties are randomly broken). Then, we feed each response’s cluster membership as a one-hot feature input to the GNN. Therefore, the GNN’s predictions are purely based on the relationships between responses in semantic space. We choose NOT to feed in the embedding vectors of responses to avoid the GNN’s dependency on textual information. This helps the GNN to generalize to questions from different domains. The overall framework is shown in Fig 1.
@@
$\ell_{q}=-\sum_{i=1}^{n}y_{i}\log\tilde{p}_{i}+(1-y_{i})\log(1-\tilde{p}_{i})$ (8)
Note that the loss is consistent marginally since the loss is minimized when $\tilde{p}_{i}=p(y_{i}|G)$. An alternative approach is to minimize the squared error $(y_{i}-\tilde{p}_{i})$, from which we get similar performances. We choose the cross-entropy loss in our work. A further consideration is to explicitly consider the similarity between $\tilde{p}_{i}$ and $\tilde{p}_{j}$ given the response similarity $w_{ij}$. We leave such exploration to the future.
@@ 3.2 Improve the estimation through multiple prompts
It is well known that the syntactic form of a question influences responses and introduces additional variance. The variance of question forms from users may pass to the variance of responses and cause even diverse responses due to the lack of confidence. To account for question variances in real applications, we analyze the LLM’s responses to multiple prompts derived from the same question. These responses are treated as answers to the same semantic question. We then apply the same method as before to predict the correctness of each response.
In particular, we rephrase the original question $q$ into $k$ different forms $\{q_{1},...,q_{k}\}$ while maintaining the original sentence’s semantic meaning. We employ a multiple rephrased questions strategy for answer sampling. Specifically, we prompt the GPT-4 to give $k$ different but with the same meaning rephrased questions for the given question $q$. Then, we sample $n/k$ responses from the LLM for each rephrased question and still get a total of $n$ responses, from which the confidence calibration is the same as we have described above. For questions about which the LLM is less certain, the model is more likely to produce diverse responses. In this scenario, confidence calibration is more accurate because the model’s uncertainty becomes more apparent.